OpenStudy (anonymous):

-12a(to the 4th) + 9a(to the 3rd) -12a(to the 2nd) = ??

OpenStudy (anonymous):

First thing you should do is factor out the equation -12a^4+9a^3-12a^2= -3a is the number you can use to factor it out. and you get -3a(4a^3+3a^2-4a)

OpenStudy (anonymous):

pull out another a

OpenStudy (anonymous):

OH sorry you can pull out -3a^2

OpenStudy (anonymous):

-3a^2 ( 4a^2 -3a +4)

OpenStudy (anonymous):

Now you can factor what's in the ( )

OpenStudy (anonymous):

So All together it goes like this (correct me if i am wrong).. -12a^4+9a^3-12a^2= -3a is the number you can use to factor it out. and you get -3a^2 ( 4a^2 -3a +4) and then you factor the inside which you get (2a+1)(2a-4) Leading to -3a^2(2a+1)(2a-4) which is the answer

OpenStudy (anonymous):

I'm getting imaginary numbers. (the square root of a negative). Is that something you've dealt with before?

OpenStudy (anonymous):

it goes like this: -3a^2(4a^2-3a+4)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You have to factor the inside (4a^2-3a+4)

OpenStudy (anonymous):

Then I used the quadratic formula to factor the inside

OpenStudy (anonymous):

what would be the result ? just multiplying and adding to simplify

OpenStudy (anonymous):

Hm Ceriona did i make a mistake?

OpenStudy (anonymous):

Oh wait (4a^2-3a+4) the +4 should be -4

OpenStudy (anonymous):

so it is (4a^2-3a-4) which leads to the answer i have put

OpenStudy (anonymous):

a = (3+ √55 i)/8 and a = (3- √55 i)/8 as imaginary roots and a=0 as a double root

OpenStudy (anonymous):

Here is what i did i dont think quadratic formula is neccesary -12a^4 +9a^3 -12a^2 -3a^2(4a^2-3a-4) -3a^2(2a-4)(2a-1) = Answer

OpenStudy (anonymous):

lol thans but u are wrong

OpenStudy (anonymous):

if there is an imaginary number ( the dicriminant is a negative number) it means that it does not pass the x axis, meaning, there is only one real root in the equation.

OpenStudy (anonymous):

There is no imaginary number though. Answer should be -3a^2(2a-4)(2a-1) if you factor it out. I dont know why she said i am wrong unless she shows me what i did wrong

OpenStudy (anonymous):

(2a-4)(2a-1)=4a^2-10a+4 not 4a^2-3a+4

OpenStudy (anonymous):

Aha thank you very much :D